OpenStudy (anonymous):

A constant torque of 19.2 N · m is applied to a grindstone for which the moment of inertia is 0.11 kg · m2. Find the angular speed after the grindstone has made 11.2 rev. Assume the grindstone starts from rest. Answer in units of rad/s

OpenStudy (anonymous):

\(\tau = I *\alpha \) \(\rightarrow \alpha = \dfrac{\tau}{I}= \dfrac{19.2}{0.11}= 174.5 ~~rad/s^2\) and you have formula \(\theta = \theta_0+ \omega_0 *t +\dfrac{1}{2}\alpha t^2\)where \(\theta = 11.2*2\pi\) so, time t = 0.898 s replace this time to formula \(\omega = \omega_0 +\alpha t\) you will have the answer is 156.7 rad/s

OpenStudy (anonymous):

Thank you! :)